OpenStudy (anonymous):

Volume of a balloon. The volume V(in cubic meters) of a hot air balloon is given by v(r)=4/3πr^3. If the radius r is increasing with time t (in seconds) according to the formula r(t)= 2/3t^3, t≥0 find the volume V as a function of the time t.

OpenStudy (ipwnbunnies):

I believe you replace r in v(r) with r(t). V(t)=4/3(pi)(2/3t^3)^3

OpenStudy (ipwnbunnies):

Expand out (2/3t^3)^3 and simplify.

OpenStudy (fibonaccichick666):

@iPwnBunnies , try asking for their ideas too. It will ensure that they are incorporated into learning the material, and they are more likely to remember it too. \[\bf \huge\color{gold}{\star}\]

OpenStudy (anonymous):

@whpalmer4 can you help?

OpenStudy (anonymous):

What do you think you should do given the provided information?

OpenStudy (anonymous):

I have no clue =(

OpenStudy (anonymous):

so far i have (2/3)^3 t^9

OpenStudy (fibonaccichick666):

Ok, so you want that in terms of t

OpenStudy (anonymous):

ok...holy crud im horrible at this stuff.. explain how please =(

OpenStudy (anonymous):

Alright. So you have r(t). In other words, the function of t. Your volume requires a radius or r. Since r is already defined, you must substitute r into the volume equation. Try it :)

OpenStudy (fibonaccichick666):

@OrthodoxMan It's the algebra, I think

OpenStudy (anonymous):

v(t)= 4/3π[(2/3)t^3]^3 = (2/3)^3 t^9 ?

OpenStudy (fibonaccichick666):

You forgot the initial constant

OpenStudy (anonymous):

Initial constant: 4/3 pi

OpenStudy (anonymous):

v= 4/3 π (8/27 t^9)

OpenStudy (fibonaccichick666):

now simplify so it isn't so ugly looking (get rid of as many fractions as you can)

OpenStudy (anonymous):

32/81πr^9

OpenStudy (anonymous):

That looks right to me.

OpenStudy (anonymous):

YAY!!!! =) thanks you guys! I dont know who to award...you all helped so much...